An IV of D5W 1,000 mL was ordered to be infused over 8 hours at a rate of 42 gtt/min. The administration set delivers 15 gtt/mL. After 4 hours, you notice only 400 mL has infused. Recalculate the gtt/min for the remaining IV solution.
Answer #1

Formula for calculating gtt/min= Volume/minutes*drop factor

Convert given hours to minutes=1hr/60min

Remaining solution is 1000ml-400ml=600ml

So volume is 600ml

Total hours is 8 hrs,but we have consumed 4 hours,hence 8hrs -4 hrs=4 hrs

Convert 4 hrs to minutes=4*60minutes=240 minutes

Drop factor=42gtt/min

600ml/240min*42gtt/min=2.5*42=105gtt/min